# Speedups in nonequilibrium thermal relaxation: Mpemba and related effects

### v1

@@ 8.2.1 Relaxation to a NESS
Motivated by water, the original formulation of the Mpemba effect is a thermal quench between Boltzmann equilibria at different temperatures. However, one can exploit the Markovian framework outlined in Sec. 2 to explore relaxations towards non-equilibrium steady states (NESS). This extension provides a first step towards the creation of more general optimal heating / cooling protocols in which the system can be driven to explore a much wider configurational space beyond those accessible when quenching only towards equilibrium steady states [65]. The boundary setup introduced in [70] naturally introduces the possibility to explore such a scenario, by coupling, e.g., different degrees of freedom to two (or more) thermal baths set at different temperatures.
@@
Figure 42: (a) Amplitude of the slowest relaxation coefficients to a NESS (Eq. 141) relaxing from an initial equilibrium at $T_{0}$ for a real (blue) and complex (green) $\lambda_{2}$. (b,c) Real and imaginary parts of $\lambda_{2}$ as a function of the two bath temperatures $T_{b_{1}}$ and $T_{b_{2}}$. (d) Phase diagram of an $N=10$ antiferromagnetic spin chain in a magnetic field $H=\sqrt{2}$ with two opposite spins coupled ($C=1$) to different thermal baths. The green area corresponds to complex-valued $\lambda_{2}$, while in the white and blue regions it is real-valued (only in the blue area an analogous of the Mpemba effect can be observed). The bisectors $T_{b_{1}}=T_{b_{2}}$ (red dashed lines) represent the equilibrium case. Figure is taken from the supplementary material of [70].
Ultimately, two possible scenarios emerge for the slowest relaxation mode (see Fig. 42): (i) $\lambda_{2}\in\mathbb{R}$, where a characterization of the Mpemba effect as for an equilibrium steady state; or (ii) $\lambda_{2}\in\mathbb{C}$ implying that $\lambda_{3}$ is its complex-conjugate and consequently the probability distribution for long enough times reads
@@ 8.2.2 Isothermal analogs
Figure 43: Particle in a periodic potential, $V(x)=V(x+L)$, subject in addition to a nonconservative driving force $f$. a) Two different initial conditions: a particle driven with a high force $f_{h}$ and a low force $f_{l}$. b) The particle driven with a high force $f_{h}$ has an initial probability distribution that is further from the equilibrium than the one driven by a low force $f_{l}$. c) After some time, the particle driven by $f_{h}$ has a probability distribution closer to equilibrium. The figure is reproduced from [219].
@@ 8.2.3 Active-particle systems
where, at each time $t$, one searches through all equilibrium (here, steady-state) distributions, as parameterized by the temperature $T_{\rm eff}$. Intuitively, one is projecting the distribution onto the closest equilibrium distribution, and associate the temperature of this equilibrium distribution to $\vec{p}(t)$. As one might expect, the effective temperature description generally breaks down for strongly nonequilibrium regimes, except some cases in mean-field systems [5]. Nevertheless, the description makes sense near equilibrium even in complicated systems. This allows the authors to study the approach to a steady state in the long-time limit. With this definition, the effective temperature can transiently dip below the bath temperature during the cooling process, so that the asymptotic dynamics has an increasing effective temperature. A similar reversal (overheating) is predicted for heating quenches and the inverse Mpemba effect, as well as in the thermodynamic limit of the mean-field antiferromagnetic Ising model [5].
Biswas and Pal have recently studied similar dynamics on discrete-state systems (three states and two velocities) [222]. They found that the eigenvalues of the rate-transition matrix of the master equation can become complex, leading to oscillations in the relaxational distance dynamics. The same authors also studied a continuous system similar to that of Ref. [221] but having an asymmetric V-shaped potential with no metastable state [223]. In that case, they found that large spontaneous velocities tend to suppress the Mpemba effect, but smaller velocities can sometimes enhance it [224].
Given the wide range of realizations of active-particle systems [220], it should be possible to test such predictions experimentally.
@@ 8.2.4 Stochastic resetting
Figure 44: a) Probability space of a three-state system. The dashed-red line shows the equilibrium locus. The relaxation with Arrhenius rates, Eq. (17), to the bath temperature, $T_{b}$, is shown with dashed-magenta (initial temperature $T_{H}$) and solid-blue lines (initial temperature $T_{C}$). The dot-dashed-green line shows the relaxation to $T_{b}$ in the presence of stochastic resetting. Here $\tau^{r}_{c}$ is the crossing time for the $L^{2}$ norm, which quantifies the onset of the Mpemba effect, and $t_{\rm SM}$ is the time at which the strong Mpemba effect occurs; i.e., $a_{2}^{(r)}(t_{\rm SM})=0$. b) Phase space of the existence of the Mpemba effect in a two-state model with stochastic resetting, with the probability of resetting to state $1$ being $\mu$, and state 2 being $1-\mu$. The energy difference $\Delta E=E_{2}-E_{1}$. The shaded area shows the region where the Mpemba effect can be observed. The figures are adapted from [228].
To appreciate how resetting affect the Mpemba effect, we recall from section 2.1.2 that the Mpemba effect does not occur for a two-state model. However, with stochastic resetting, one can also have a Mpemba effect in the two-state system. Moreover, it is enough to consider resetting to a single state. Generalizing to resetting to both states and introducing the probability of resetting to state 1 is $\Delta_{1}=\mu$ and to state 2 is $\Delta_{2}=1-\mu$ where $\mu\in[0,1]$ we have the following condition for the occurrence of the Mpemba effect
$\frac{\Delta E}{T_{b}}<\ln\frac{\mu}{1-\mu},$ (149)
where $\Delta E\equiv E_{2}-E_{1}$. The phase space $(\mu,\Delta E/T_{b})$ is shown on Fig. 44b.
@@
Figure 45: Pareto front showing the optimal protocol that minimizes the crossing time $\tau_{\rm c}^{(r)}$ and the dissipation to the environment, $\Delta S^{(r)}_{\rm env}$ for a fully connected four-state Markov system. The solid blue line represents resetting to a single state, while the dashed-black line represents resetting to two states (1 and 4 here) with $\Delta=\mu\delta_{i,1}+(1-\mu)\delta_{i,4}$. Suboptimal choices of resetting probability (to state 1) $\mu$ are above the dashed black line. For protocols in the shaded area, both the crossing time $\tau_{c}^{(r)}$ and the dissipation to the environment $\Delta S^{(r)}_{\rm en}$ can be simultaneously optimized, hence corresponding to a single optimal point $r^{\rm opt}=r_{\rm min}$. The figure is adapted from [228].
@@
Additional connection between stochastic resetting and Mpemba were discussed in  [200], where stochastic resetting was utilized to optimize the relaxation process of an autonomous Maxwell’s demon by inducing a strong Mpemba effect. Maxwell’s demon is introduced and discussed in Sec. 8.1.2. Specifically, the manuscript us the exactly solvable demon model proposed in [206]. By introducing a discrete-time stochastic resetting mechanism and applying it to the demon setup, the authors showed that the time for the autonomous demon to relax to its periodic steady state can be reduced by a protocol inspired by a strong Mpemba effect. The design strategy is illustrated for the two-state Maxwell demon shown in Fig. 46a. This model system consists of a demon with two states and an infinitely long tape – a stream of information with bits 00 and $1$. The combined system is shown on Fig. 46b. The demon starts the cycle at a temperature of one of the reservoirs, and, during a time $\tau$, interacts with the memory tape. There is no friction, and the demon interacts with only a single bit at a time. During the interaction, the demon undergoes an intrinsic transition between its pairs of states. The bit can also change its state. The two transitions can occur simultaneously and lead to anomalous work production. The demon’s disordered transitions can be rectified by incoming bits. The combined system of the demon and the tape obeys a Master equation. After a certain time, it reaches a periodic steady state.
@@
Figure 46: a) Two-state Maxwell demon interacting with a tape (sequence of bits) and two reservoirs at temperatures $T_{h}$ and $T_{c}$. b) A composite four-state system consisting of demon and tape. The black edges correspond to transitions at high temperature $T_{h}$, and the red edge corresponds to the transition at low temperature $T_{c}$. The figures are adapted from [200].
@@
The authors introduce a strategy of resetting the demon with a constant probability in discrete time for a fixed amount of time $t_{c}=N_{c}\tau$, such that the resulting evolution does not have the coefficient to the slowest mode, that is $a_{2}^{(r)}(N_{c})=0$. They show that the sufficient condition to find a physical critical number of periods, $N_{c}\geq 1$ is $a_{2}/d_{2}\leq 0$, where $d_{2}$, as before, is the coefficient of the resetting state $\vec{\Delta}$ in front of the second right eigenvector $\vec{v}_{2}$ and $a_{2}$ is the coefficient to the second eigenvector in the absence of resetting. This is in agreement with the above-described result appeared in [228] c.f. Eq. (148). The authors show how stochastic resetting can notably enhance the performance of an autonomous Maxwell’s demon in two ways: the speed of reaching their functional state and the range of their effective work regions. Namely, they show that one can drive the autonomous demon system to its functional periodic steady state with a shortened relaxation time by resetting the demon at a predetermined critical time and then stopping the resets. They also show how, by continuing with the resets, one can increase the effective working area of Maxwell’s demon. Lastly, they discover a dual-functioning region of the demon with resetting, where it can both produce work and erase information at the same time. The dual region comes at the cost of resetting.
@@
Lastly, a Maxwell demon model interacting with a tape in connection with the Mpemba effect was discussed in [98]. Similar to previous research, the authors identify conditions under which the relaxation to the functioning state is accelerated. In [98], the tape is maintained at a finite temperature, and the dynamics of relaxation are explored as a function of a control parameter (load factor, see also Sec. 8.1.4). We reference this work in the context of optimizing the performance of the device in Sec. 8.1.2.
@@ 8.2.6 Controlled state transformations
Since simple protocols such as steps in temperature or some driving force lead to interesting results, it is natural to consider more general protocols $\vec{\theta}(t)$ that may include multiple (even an infinite number) of parameters and may or may not involve temperature. This is the domain of control theory, which is the subject of increasing interest among physicists [196]. In the framework of control theory, these protocols all aim to implement state-to-state transitions. The protocols are typically finite-time protocols, lasting a time $\tau$. The states either reach the desired state at the end of the protocol or reach it after a passive relaxation that starts at time $t=\tau$. For the latter case, one can still define a notion of protocol relaxation time using the magnitude of the $a_{2}$ coefficient, as discussed in Sec. 2.1.1.
In general, there are many protocols $\vec{\theta}(t)$—often, infinitely many—that transform a given state into a specified final state. Sometimes finding a single feasible protocol, a $\vec{\theta}(t)$ that takes a system from its initial state $x(0)$ to a desired final state $x(\tau)$, is enough. Though usually the transformation should satisfy additional considerations. One that is nearly always relevant is that the allowed values of a control parameter are typically limited: for example, applied forces must be below some value. Beyond constraints on the values of a control parameter, one might seek to make the transformation in minimum time, to minimize dissipation, or to follow some prescribed path.
@@
More recently, Patrón et al. studied optimal bang-bang protocols in harmonic potentials [243], which they referred to as leading to the “thermal brachistochrone,” (minimum-time trajectory). The term refers to the classic problem posed by Johann Bernoulli and solved by Isaac Newton of the shape with the fastest transition time for a ballistic particle [244]. For harmonic potentials, the nonequilibrium distributions are always Gaussian at all times in the protocol, which allows for transitions where the equilibrium is reached in finite time. This is generally not possible for non-Gaussian potentials when controlling only the temperature. More recently, the same authors have generalized their calculation to the nonequilibrium steady states created in a Brownian gyrator, where a particle trapped in a two-dimensional harmonic oscillator is subject to different noise strengths along different coordinate axes and rotates as a result [245].
Finally, we discuss a more sophisticated control extension of the Mpemba effect proposed by Chittari and Lu [246]. They seek to establish conditions for the existence of rapid (“hasty”) protocols that speed up operations that transform a system from some initial state to some desired end state in a finite time. The basic idea is to manipulate the location of manifolds describing fast and slow relaxations such that a trajectory can always proceed along fast segments; hence, the term “hasty.” Chittari and Lu argue that Mpemba effect-like shortcuts only constitute a small fraction of the diverse categories of hasty shortcuts and refer to the Mpemba effect as a “one-step” hasty shortcut since the system is “steered away” from slow dynamics via a single control step (by setting the initial condition). Moreover, that single control step is what the authors call “counter-shooting only” as the controls go beyond the range of the initial and final (i.e., starting from a hotter initial temperature to facilitate faster cooling to the environment).

### v3

@@ 8.2.1 Relaxation to a NESS
Motivated by water, the original formulation of the Mpemba effect is a thermal quench between Boltzmann equilibria at different temperatures. However, one can exploit the Markovian framework outlined in Sec. 2 to explore relaxations towards nonequilibrium steady states (NESS). This extension provides a first step towards the creation of more general optimal heating / cooling protocols, in which the system can be driven to explore a much wider configurational space beyond those accessible when quenching only towards equilibrium steady states [65]. The boundary setup introduced in [70] naturally introduces the possibility to explore such a scenario, by coupling, e.g., different degrees of freedom to two (or more) thermal baths set at different temperatures.
@@
Figure 42: (a) Amplitude of the slowest relaxation coefficients to a NESS (Eq. 141) relaxing from an initial equilibrium at $T_{0}$ for a real (blue) and complex (green) $\lambda_{2}$. (b,c) Real and imaginary parts of $\lambda_{2}$ as a function of the two bath temperatures $T_{b_{1}}$ and $T_{b_{2}}$. (d) Phase diagram of an $N=10$ antiferromagnetic spin chain in a magnetic field $H=\sqrt{2}$ with two opposite spins coupled ($C=1$) to different thermal baths. The green area corresponds to complex-valued $\lambda_{2}$, while in the white and blue regions it is real-valued (only in the blue area an analogous of the Mpemba effect can be observed). The bisectors $T_{b_{1}}=T_{b_{2}}$ (red dashed lines) represent the equilibrium case. Source: Reprinted with permission from the supplementary material of [70].
Ultimately, two possible scenarios emerge for the slowest relaxation mode; see Fig. 42: (i) $\lambda_{2}\in\mathbb{R}$, where a characterization of the Mpemba effect as for an equilibrium steady state; or (ii) $\lambda_{2}\in\mathbb{C}$ implying that $\lambda_{3}$ is its complex-conjugate, and consequently the probability distribution for long enough times is
@@ 8.2.2 Isothermal analogs
Figure 43: Particle in a periodic potential, $V(x)=V(x+L)$, subject in addition to a nonconservative driving force $f$. (a) Two different initial conditions: a particle driven with a high force $f_{h}$ and a low force $f_{l}$. (b) The particle driven with a high force $f_{h}$ has an initial probability distribution that is further from the equilibrium than the one driven by a low force $f_{l}$. (c) After some time, the particle driven by $f_{h}$ has a probability distribution closer to equilibrium. Source: Reprinted with permission from [229].
@@ 8.2.3 Active-particle systems
where, at each time $t$, one searches through all equilibrium (here, steady-state) distributions, as parameterized by the temperature $T_{\rm eff}$. Intuitively, one is projecting the distribution onto the closest equilibrium distribution, and associate the temperature of this equilibrium distribution to $\vec{p}(t)$. As one might expect, the effective temperature description generally breaks down for strongly nonequilibrium regimes, except in some cases in mean-field systems [5]. Nevertheless, the description makes sense near equilibrium, even in complicated systems. This allows the authors to study the approach to a steady state in the long-time limit. With this definition, the effective temperature can transiently dip below the bath temperature during the cooling process, so that the asymptotic dynamics has an increasing effective temperature. A similar reversal (overheating) is predicted for heating quenches and the inverse Mpemba effect, as well as in the thermodynamic limit of the mean-field antiferromagnetic Ising model [5].
Biswas and Pal have recently studied similar dynamics on discrete-state systems (three states and two velocities) [232]. They found that the eigenvalues of the rate-transition matrix of the master equation can become complex, leading to oscillations in the relaxational distance dynamics. The same authors also studied a continuous system similar to that of Ref. [231] but having an asymmetric V-shaped potential with no metastable state [71]. In that case, they found that large spontaneous velocities tend to suppress the Mpemba effect, but smaller velocities can sometimes enhance it [233]. Given the wide range of realizations of active-particle systems [230], it should be possible to test such predictions experimentally.
@@ 8.2.4 Stochastic resetting
Figure 44: (a) Probability space of a three-state system. The dashed-red line shows the equilibrium locus. The relaxation with Arrhenius rates, Eq. (17), to the bath temperature, $T_{b}$, is shown with dashed-magenta (initial temperature $T_{H}$) and solid-blue lines (initial temperature $T_{C}$). The dot-dashed-green line shows the relaxation to $T_{b}$ in the presence of stochastic resetting. Here $\tau^{r}_{c}$ is the crossing time for the $L^{2}$ norm, which quantifies the onset of the Mpemba effect, and $t_{\rm SM}$ is the time at which the strong Mpemba effect occurs; i.e., $a_{2}^{(r)}(t_{\rm SM})=0$. (b) Phase space of the existence of the Mpemba effect in a two-state model with stochastic resetting, with the probability of resetting to State $1$ being $\mu$, and State 2 being $1-\mu$. The energy difference $\Delta E=E_{2}-E_{1}$. The shaded area shows the region where the Mpemba effect can be observed. Source: Adapted with permission from [237].
To appreciate how resetting affect the Mpemba effect, we recall from Sec. 2.1.2 that the Mpemba effect does not occur for a two-state model. However, with stochastic resetting, one can also have a Mpemba effect in the two-state system. Moreover, it is enough to consider resetting to a single state. Generalizing to resetting to both states and introducing the probability of resetting to State 1 is $\Delta_{1}=\mu$ and to State 2 is $\Delta_{2}=1-\mu$ where $\mu\in[0,1]$ we have the following condition for the occurrence of the Mpemba effect
$\frac{\Delta E}{k_{B}T_{b}}<\ln\frac{\mu}{1-\mu},$ (149)
where $\Delta E\equiv E_{2}-E_{1}$. The phase space $(\mu,\Delta E/k_{B}T_{b})$ is shown on Fig. 44b.
@@
Figure 45: Pareto front showing the optimal protocol that minimizes the crossing time $\tau_{\rm c}^{(r)}$ and the dissipation to the environment, $\Delta S^{(r)}_{\rm env}$ for a fully connected four-state Markov system. The solid blue line represents resetting to a single state, while the dashed-black line represents resetting to two states (1 and 4 here) with $\Delta=\mu\delta_{i,1}+(1-\mu)\delta_{i,4}$. Suboptimal choices of resetting probability (to state 1) $\mu$ are above the dashed black line. For protocols in the shaded area, both the crossing time $\tau_{c}^{(r)}$ and the dissipation to the environment $\Delta S^{(r)}_{\rm en}$ can be simultaneously optimized, hence corresponding to a single optimal point $r^{\rm opt}=r_{\rm min}$. Source: Adapted with permission from [237].
@@
Additional connection between stochastic resetting and Mpemba were discussed in  [209], where stochastic resetting was utilized to optimize the relaxation process of an autonomous Maxwell’s demon by inducing a strong Mpemba effect. Maxwell’s demon is introduced and discussed in Sec. 8.1.2. Specifically, the manuscript us the exactly solvable demon model proposed in [215]. By introducing a discrete-time stochastic resetting mechanism and applying it to the demon setup, the authors showed that the time for the autonomous demon to relax to its periodic steady state can be reduced by a protocol inspired by a strong Mpemba effect. The design strategy is illustrated for the two-state Maxwell demon shown in Fig. 46a. This model system consists of a demon with two states and an infinitely long tape, a stream of information with bits 00 and $1$. The combined system is shown on Fig. 46b. The demon starts the cycle at a temperature of one of the reservoirs, and, during a time $\tau$, interacts with the memory tape. There is no friction, and the demon interacts with only a single bit at a time. During the interaction, the demon undergoes an intrinsic transition between its pairs of states. The bit can also change its state. The two transitions can occur simultaneously and lead to anomalous work production. The demon’s disordered transitions can be rectified by incoming bits. The combined system of the demon and the tape obeys a master equation. After a certain time, it reaches a periodic steady state.
@@
Figure 46: (a) Two-state Maxwell demon interacting with a tape (sequence of bits) and two reservoirs at temperatures $T_{h}$ and $T_{c}$. (b) A composite four-state system consisting of demon and tape. The black edges correspond to transitions at high temperature $T_{h}$, and the red edge corresponds to the transition at low temperature $T_{c}$. Source: Adapted with permission from [209].
@@
The authors introduce a strategy of resetting the demon with a constant probability in discrete time for a fixed amount of time $t_{c}=N_{c}\tau$, such that the resulting evolution does not have the coefficient to the slowest mode, that is $a_{2}^{(r)}(N_{c})=0$. They show that the sufficient condition to find a physical critical number of periods, $N_{c}\geq 1$ is $a_{2}/d_{2}\leq 0$, where $d_{2}$, as before, is the coefficient of the resetting state $\vec{\Delta}$ in front of the second right eigenvector $\vec{v}_{2}$ and $a_{2}$ is the coefficient to the second eigenvector in the absence of resetting. This agrees with the above-described result from [237] c.f. Eq. (148). The authors show how stochastic resetting can notably enhance the performance of an autonomous Maxwell’s demon in two ways: the speed of reaching their functional state and the range of their effective work regions. Namely, they show that one can drive the autonomous demon system to its functional periodic steady state with a shortened relaxation time by resetting the demon at a predetermined critical time and then stopping the resets. They also show how, by continuing with the resets, one can increase the effective working area of Maxwell’s demon. Lastly, they discover a dual-functioning region of the demon with resetting, where it can both produce work and erase information at the same time. The dual region comes at the cost of resetting.
@@
A Maxwell demon model interacting with a tape in connection with the Mpemba effect was discussed in [97]. Similar to previous research, the authors identify conditions under which the relaxation to the functioning state is accelerated. In [97], the tape is maintained at a finite temperature, and the dynamics of relaxation are explored as a function of a control parameter (load factor, see also Sec. 8.1.4). We reference this work in the context of optimizing the performance of the device in Sec. 8.1.2. Finally, for quantum systems, the acceleration of relaxation in Markovian open quantum systems through quantum reset processes was considered in  [238].
@@ 8.2.6 Controlled state transformations
Since simple protocols such as steps in temperature or some driving force lead to interesting results, it is natural to consider more general protocols $\vec{\theta}(t)$ that may include multiple (even an infinite number) of parameters and may or may not involve temperature. This is the domain of control theory, a subject of increasing interest to physicists [205]. In the framework of control theory, these protocols all aim to implement state-to-state transitions. The protocols are typically finite-time protocols, lasting a time $\tau$. The states either reach the desired state at the end of the protocol or reach it after a passive relaxation that starts at time $t=\tau$. For the latter case, one can still define a notion of protocol relaxation time using the magnitude of the $a_{2}$ coefficient, as discussed in Sec. 2.1.1.
In general, there are many protocols $\vec{\theta}(t)$—often, infinitely many—that transform a given state into a specified final state. Sometimes finding a single feasible protocol, a $\vec{\theta}(t)$ that takes a system from its initial state $x(0)$ to a desired final state $x(\tau)$, is enough. Usually, though, the transformation should satisfy additional considerations. One that is nearly always relevant is that the allowed values of a control parameter are typically limited: for example, applied forces must be below some value. Beyond constraints on the values of a control parameter, one might seek to make the transformation in minimum time, to minimize dissipation, or to follow some prescribed path.
@@
More recently, Patrón et al. studied optimal bang-bang protocols in harmonic potentials [253], which they referred to as leading to the “thermal brachistochrone,” (minimum-time trajectory). The term refers to the classic problem posed by Johann Bernoulli and solved by Isaac Newton of the shape with the fastest transition time for a ballistic particle [254]. For harmonic potentials, the nonequilibrium distributions are always Gaussian at all times in the protocol, which allows for transitions where the equilibrium is reached in finite time. This is generally not possible for non-Gaussian potentials when controlling only the temperature. Furthermore, the same authors have generalized their calculation to the nonequilibrium steady states created in a Brownian gyrator, where a particle trapped in a two-dimensional harmonic oscillator is subject to different noise strengths along different coordinate axes and rotates as a result [255]. In [256] dynamics invariants, based on shortcuts to adiabaticity, are used to speedup the equilibration of an open quantum system. The speedup is compared to that resulting from a Mpemba effect.
Finally, we discuss a more sophisticated control extension of the Mpemba effect proposed by Chittari and Lu [257]. They seek to establish conditions for the existence of rapid (“hasty”) protocols that speed up operations that transform a system from some initial state to some desired end state in a finite time. The basic idea is to manipulate the location of manifolds describing fast and slow relaxations such that a trajectory can always proceed along fast segments; hence, the term “hasty.” Chittari and Lu argue that Mpemba effect-like shortcuts only constitute a small fraction of the diverse categories of hasty shortcuts and refer to the Mpemba effect as a “one-step” hasty shortcut since the system is “steered away” from slow dynamics via a single control step (by setting the initial condition). Moreover, that single control step is what the authors call “counter-shooting only,” as the controls go beyond the range of the initial and final bounds. That is, they start from a hotter initial temperature to facilitate faster cooling to the environment.